# Test-Time Scaling with Reflective Generative Model

### v1

@@ 4.1 Baseline&Dataset
We conduct experiments on the models with three different sizes: MetaStone-S1-1.5B, 7B, and 32B, which are initialized from DeepSeek-R1-Distill-Qwen-1.5B/7B (Guo et al., 2025), and QWQ-32B (Team, 2025) with continual reinforcement training. Our training dataset is constructed from multiple publicly available math-related sources, including NuminaMath (Li et al., 2024), OpenR1-Math-220k, DeepScaleR (Luo et al., 2025), LIMR (Li et al., 2025), and OREAL-RL (Lyu et al., 2025). We apply a multi-agent data cleaning framework to ensure data quality, resulting in a final dataset of 40k high-quality examples. In the inference stage, we set 3 reasoning efforts with $k=2,8,32$ in Eq.3, named MetaStone-S1-low, -medium, and -high.
@@
Small-size Open-Sourced Models
@@
Large-size Open-Sourced Models
@@
Closed-Sourced Models
@@
OpenAI-o3-mini* 79.6 74.8 67.4 75.9
@@
Table 1: Comparison of MetaStone-S1 models and other comparable models. * denotes medium level of OpenAI-o3-mini. The best and second-best results are shown in bold and underlined.
@@ 4.2 Main Results
Table 1 summarizes the performance of our MetaStone-S1 models across four representative benchmarks. We denote baseline as the models without SPRM. Across different model scales, our proposed SPRM strategy consistently enhances the baseline, particularly on mathematical reasoning benchmarks. Specifically, compared with the baseline, MetaStone-S1 achieves performance gains of 18.6/15.5/5.3 points on AIME24 and 10.5/7.4/3.1 points on AIME25 for the 1.5B/7B/32B sizes, respectively. For other tasks, SPRM continues to offer stable improvements, yielding gains of 5.7/5.0/0.8 points on LiveCodeBench and 2.3/6.5/0.3 points on C-EVAL.
@@
We further compare MetaStone-S1 against both advanced open-source and closed-source models. Among the open-source models, we consider DeepScaleR-1.5B-Preview (Luo et al., 2025), DeepSeek-R1-Distill-Qwen (1.5B/7B/32B), DeepSeek-R1-Distill-LLaMA-8B (Guo et al., 2025), QwQ-32B (Team, 2025), GLM-Z1-32B-0414 (GLM et al., 2024), s1-32B (Muennighoff et al., 2025), and DeepSeek-R1-671B (Guo et al., 2025). Among the closed-source models, we include Claude-3.5-Sonnet-1022, GPT-4o-0522, and OpenAI-o3-mini-medium.
@@
At the small scale, MetaStone-S1-1.5B and MetaStone-S1-7B consistently outperform the listed open-sourced models with comparable or larger parameter sizes. Especially, MetaStone-S1-1.5B-low surpasses DeepScaleR-1.5B-Preview and R1-Distill-Qwen-1.5B on all datasets. And MetaStone-S1-1.5B-high further outperforms R1-Distill-Qwen-7B and R1-Distill-Llama-8B on AIME24 (57.9% vs 55.5%), demonstrating strong efficiency and capability of our lightweight SPRM. For MetaStone-S1-7B, its low reasoning effort has outperformed R1-Distill-Qwen-7B and R1-Distill-Llama-8B on AIME24 and LiveCodeBench. And MetaStone-S1-7B-high further gains the improvement of 14.7 points on AIME24 (70.2% vs 55.5%) and 4.8 points on LiveCodeBench (44.4% vs 39.6%).
@@
At the larger scale, MetaStone-S1-32B-high achieves superior results on mathematical reasoning tasks, outperforming all listed open-sourced models of comparable or even larger size by +4.4% on AIME24 (85.2% vs 80.8%) and +3.6% on AIME25 (73.6% vs 70.0%). For other out-of-distribution tasks, MetaStone-S1-32B-high surpasses other 32B-sized models by +0.8% on LiveCodeBench (63.4% vs 64.2%) and +1.3% on C-EVAL (89.7% vs 88.4%). Compared to closed-source models, medium and high levels of MetaStone-S1-32B outperform Claude-3.5-Sonnet-1022 and GPT-4o-0522, and achieving comparable performance with medium level of OpenAI-o3-mini (85.2% vs 79.6% on AIME24, 73.6% vs 74.8% on AIME25, 64.2% vs 67.4% on LiveCodeBench, 89.7% vs 75.9% on C-EVAL,), highlighting its strong competitiveness in both mathematical and general reasoning tasks.
4.3 Ablation Study
Effectiveness of SPRM.
As shown in Table.2, our SPRM, with only around 26M extra parameters, achieves higher performance compared with 72B ORM and PRM. This demonstrates that sharing parameters between the reward model and policy model is able to generate high-quality guidance without the requirement of additional large-scale reward models. To further analyze the efficacy of different reward models, we compare their performance under different candidates $k$ in Fig.3. Across different $k$ and model sizes, SPRM consistently outperforms other methods, indicating its strong ability to distinguish between high and low quality reasoning trajectories.
To better show the effect of SPRM, in Fig.4, we visualize an example. The proposed SPRM correctly detect the calculation error at corresponding steps, results in a low score on both step and sample level.
@@ 4.3 Ablation Study
Figure 3: Evaluation of different number of candidates on AIME24.
[…]
Figure 4: SPRM predcitions on reasoning trajectory. Only key steps are listed. Correct and error steps are marked in green and red.
@@
We evaluate the effectiveness of SPRLoss in Table.3. Compared with using the final answer correctness as step-level supervision for PRM training, our proposed self-supervised optimization method achieves larger performance gains on both 1.5B and 7B models. Furthermore, Fig.5 shows the prediction score gap between correct and incorrect solutions. Compared to the BCELoss, SPRLoss demonstrates stronger discriminative capability with a larger score gap. This indicates that treating final answer correctness as step-level labels introduces substantial label noise, which harms the optimization. In contrast, SPRLoss leverages self-supervised signals to reduce the impact of noisy supervision, leading to more stable and accurate training.
@@ 4.4 Scaling Law of the Reflective Generative Model
Figure 6: The scaling law of Reflective Generative Models.
4.5 Aha Moment of SPRM
In Fig. 7, we present the predicted scores of SPRM for both correct and incorrect reasoning trajectories throughout the training process. During the initial phase of training, the optimization trajectories for all samples follow a similar trend, indicating that SPRM has not yet learned to distinguish between correct and incorrect reasoning trajectories. However, after a certain number of training steps (e.g., around 10/20/50 steps, 1280/2560/6400 samples for MetaStone-S1-1.5B/7B/32B, respectively), we observe a distinct "aha moment" point where the optimization trends of different reasoning trajectories begin to diverge. This suggests that the model is starting to judge the correctness based on the reasoning contents. With this aha moment, SPRM can be progressively refined through our proposed self-supervised SPRLoss, leading to a clear score gap between correct and incorrect reasoning trajectories, and enabling effective TTS.
Specially, we show an example in Fig.8. We fix the reasoning trajectory and use SPRM with different steps for scoring. In this case, the model mistakenly confuses $\varepsilon_{k}$ and $\varepsilon_{k+1}$, resulting in an incorrect solution. The SPRM fails to recognize the error before the aha moment, while the SPRM after the aha moment can correctly discriminate it.
Figure 7: The training process of SPRM. The blue and red curves denote the final prediction score on correct and incorrect reasoning trajectories. The green dashed line indicates the "aha moment".
Example: Aha Moment of SPRM Question: Let the sequence of rationals $x_{1},x_{2},\dots$ be defined such that $x_{1}=\frac{25}{11}$ and $[x_{k+1}=\frac{1}{3}\left(x_{k}+\frac{1}{x_{k}}-1\right)].$ $x_{2025}$ can be expressed as $\frac{m}{n}$ for relatively prime positive integers $m$ and $n$. Find the remainder when $m+n$ is divided by $1000$. Reasoning Trajectories: Alternatively, perhaps make a substitution. Let me define $\mu_{k}=1/\varepsilon_{k}$. Then, $1/\varepsilon_{k+1}=-1/\varepsilon_{k}+(8/9)/\varepsilon_{k}²$ Multiply both sides by $\varepsilon_{k}²$: $\varepsilon_{k}=-\varepsilon_{k}+(8/9)$ Thus, $\varepsilon_{k}+\varepsilon_{k+1}=8/9$ But no, not quitequestion and the correct answer should be 21+49=70) Before aha moment: Step score: 0.52; Final score: 0.54 After aha moment: Step score: 0.12; Final score: 0.36
Figure 8: Comparision of PRM predcitions before and after aha moment. Only key steps are listed. The error steps are marked in red.
4.6 Extend on MCTS
Since SPRM produces step-level evaluations, it can be naturally used in step-level search-based TTS methods such as Monte Carlo Tree Search (MCTS). In our setup, instead of performing full simulations to the end of a solution trajectory, we directly use SPRM to estimate the value of each node during the search. This enables a more efficient evaluation at each expansion step. In the expanding stage, we expand 4 children for the selected node and generate 1024 tokens in each child node. To balance the computation cost, we set the maximum number of tokens in the MCTS process from 0 (without MCTS) to 160k for each question.
Table.4 presents the performance of MCTS on the AIME24 dataset. By increasing the maximum number of searching tokens, the performance improved from 39.3 to 52.8, demonstrating the effectiveness of SPRM in providing high-quality step-level guidance in the reasoning stage. However, the performance of MCTS is still lower than the results under the Best-of-N strategy reported in Table 1. This is primarily due to the computational overhead inherent in tree-based search methods, which leads to incomplete searching in our setting. Nonetheless, the observed gains over the baseline validate the potential of our SPRM for integration with more advanced search-based inference methods.
@@ 5 Conclusion
In this work, we introduce MetaStone-S1, a novel Reflective Generative model that achieves competitive performance with OpenAI’s o3 model. In MetaStone-S1, we propose a self-supervised process reward model (SPRM) which provides fine-grained step-level guidance without requiring manual annotations. SPRM introduces only a lightweight prediction head and shares the majority of its parameters with the policy model, allowing on-policy joint optimization. The Reflective Generative models with SPRM naturally suitable for test time scaling, enabling performance improvement across both in-domain and out-of-domain benchmarks. We also summarized a scaling law between total thinking length and TTS performance. Furthermore, we find that our SPRM exhibits strong potential when integrated with search-based TTS methods, such as Monte Carlo Tree Search (MCTS). In future work, we will further explore its capability on step-level searching-based TTS and investigate its potential to enable real-time reasoning enhancement during inference.

### v2

@@ 5.1 Baseline&Dataset
We conduct experiments on the models with three different sizes: MetaStone-S1-1.5B, 7B, and 32B, which are initialized from DeepSeek-R1-Distill-Qwen-1.5B/7B (Guo et al., 2025), and QWQ-32B (Team, 2025) with continual reinforcement training. After adding the SPRM head, only 5M/26M/53M extra parameters are introduced for MetaStone-S1-1.5B/7B/32B. Our training dataset is constructed from multiple publicly available math-related sources, including NuminaMath (Li et al., 2024), OpenR1-Math-220k, DeepScaleR (Luo et al., 2025), LIMR (Li et al., 2025), and OREAL-RL (Lyu et al., 2025). We apply a multi-agent data cleaning framework to ensure data quality, resulting in a final dataset of 40k high-quality examples. In the inference stage, we set 3 reasoning efforts with $k=2,8,32$ in Eq.7, named MetaStone-S1-low, -medium, and -high.
@@
Small-size Open-Source Models
@@
Large-size Open-Source Models
@@
Closed-Source Models
@@
OpenAI o1-mini 63.6 50.7 53.8 68.9
OpenAI o1-1217 79.2 - 63.4 -
OpenAI o3-mini* 79.6 74.8 67.4 75.9
@@
Table 1: Comparison of MetaStone-S1 models and other comparable models. * denotes medium level of OpenAI o3-mini. The best and second-best results are shown in bold and underlined.
@@ 5.2 Main Results
Table 1 summarizes the performance of our MetaStone-S1 models across four representative benchmarks. We denote baseline as the only policy model without using the Reflective Generative Form. Across different model scales, our proposed Reflective Generative Form consistently enhances the baseline, particularly on mathematical reasoning benchmarks. Specifically, compared with the baseline, MetaStone-S1 achieves performance gains of 18.6/15.5/5.3 points on AIME24 and 10.5/7.4/3.1 points on AIME25 for the 1.5B/7B/32B sizes, respectively. For other tasks, our Reflective Generative Form continues to offer stable improvements, yielding gains of 5.7/5.0/0.8 points on LiveCodeBench and 2.3/6.5/0.3 points on C-Eval.
@@
We further compare MetaStone-S1 against both advanced open-source and closed-source models. Among the open-source models, we consider DeepScaleR-1.5B-Preview (Luo et al., 2025), DeepSeek-R1-Distill-Qwen (1.5B/7B/32B), DeepSeek-R1-Distill-LLaMA-8B (Guo et al., 2025), QwQ-32B (Team, 2025), GLM-Z1-32B-0414 (GLM et al., 2024), s1-32B (Muennighoff et al., 2025), and DeepSeek-R1-671B (Guo et al., 2025). Among the closed-source models, we include Claude-3.5-Sonnet-1022, GPT-4o-0522, OpenAI o1-mini, OpenAI o1-1217, and OpenAI o3-mini-medium.
@@
At the small scale, MetaStone-S1-1.5B and MetaStone-S1-7B consistently outperform the listed open-source models with comparable or larger parameter sizes. Especially, MetaStone-S1-1.5B-low surpasses DeepScaleR-1.5B-Preview and R1-Distill-Qwen-1.5B on all datasets. And MetaStone-S1-1.5B-high further outperforms R1-Distill-Qwen-7B and R1-Distill-Llama-8B on AIME24 (57.9% vs 55.5%), demonstrating strong efficiency and capability of our lightweight SPRM. For MetaStone-S1-7B, its low reasoning effort has outperformed R1-Distill-Qwen-7B and R1-Distill-Llama-8B on AIME24 and LiveCodeBench. And MetaStone-S1-7B-high further gains the improvement of 14.7 points on AIME24 (70.2% vs 55.5%) and 4.8 points on LiveCodeBench (44.4% vs 39.6%).
@@
At the larger scale, MetaStone-S1-32B-high achieves superior results on mathematical reasoning tasks, outperforming all listed open-source models of comparable or even larger size by +4.4% on AIME24 (85.2% vs 80.8%) and +3.6% on AIME25 (73.6% vs 70.0%). For other out-of-distribution tasks, MetaStone-S1-32B-high surpasses other 32B-sized models by +0.8% on LiveCodeBench (64.2% vs 63.4%) and +1.3% on C-Eval (89.7% vs 88.4%). Compared to closed-source models, medium and high levels of MetaStone-S1-32B outperform Claude-3.5-Sonnet-1022 and GPT-4o-0522, and achieving comparable performance with medium level of OpenAI o3-mini (85.2% vs 79.6% on AIME24, 73.6% vs 74.8% on AIME25, 64.2% vs 67.4% on LiveCodeBench, 89.7% vs 75.9% on C-Eval,), highlighting its strong competitiveness in both mathematical and general reasoning tasks.
Figure 3: The scaling law of Reflective Generative Models.
5.3 Scaling Law of the Reflective Generative Model
Figure 4: The training process of SPRM. The blue and red curves denote the final score on correct and incorrect reasoning trajectories. The green dashed line indicates the "aha moment".
Figure 5: Comparison of SPRM’s predictions before and after the aha moment. Only key steps are listed. The error steps are marked in red.
5.4 Aha Moment of the Reflective Generative Model
In Fig. 4, we present the final evaluation scores from MetaStone-S1 for both correct and incorrect reasoning trajectories throughout the training process. During the initial phase of training, the optimization trajectories for all samples follow a similar trend, indicating that the model has not yet learned to distinguish between correct and incorrect reasoning trajectories. However, after a certain number of training steps (e.g., around 10/20/50 steps, 1280/2560/6400 samples for MetaStone-S1-1.5B/7B/32B, respectively), we observe a distinct "aha moment" point where the optimization trends of different reasoning trajectories begin to diverge. This suggests that the model is starting to judge the correctness based on the reasoning contents. With this aha moment, MetaStone-S1 can progressively refine its SPRM head through our proposed self-supervised SPRLoss, leading to a clear score gap between correct and incorrect reasoning trajectories, and enabling effective TTS.
Specially, we show an example in Fig.5. We fix the reasoning trajectory and use MetaStone-S1 before and after the aha moment for scoring. In this case, the model mistakenly confuses $\varepsilon_{k}$ and $\varepsilon_{k+1}$, resulting in an incorrect solution. Our model fails to recognize the error before the aha moment, while the model after the aha moment can correctly discriminate it.
@@ 5.4 Aha Moment of the Reflective Generative Model
Figure 6: Evaluation of varying numbers of candidate reasoning trajectories on AIME24.
Figure 7: SPRM’s predictions on reasoning trajectories. Only key steps are listed. Correct and error steps are marked in green and red.
5.5 Ablation Study
Effectiveness of SPRM.
As shown in Table.2, our SPRM, with only around 26M extra parameters, achieves higher performance compared with 72B ORM and PRM. This demonstrates that sharing parameters between the reward model and policy model is able to generate high-quality guidance without the requirement of additional large-scale reward models. To further analyze the efficacy of different reward models, we compare their performance under different numbers of candidate reasoning trajectories $k$ in Fig.6. Across different $k$ and model sizes, SPRM consistently outperforms other methods, indicating its strong ability to distinguish between high and low quality reasoning trajectories.
Fig.7 shows the visualization of step-wise evaluation scores from SPRM. It can be observed that SPRM effectively identifies low-quality processes generated by the policy model, including logical error (e.g. the misunderstanding of b in step 58 of example 1) and calculation error (e.g., the incorrect computation 9 × 21 + 7 = 193 in step 32 of example 2). SPRM assigns low scores to these low-quality steps. Since SPRM only outputs process scores, we additionally provide post-descriptions within the dashed boxes for better clarity.
@@ 5.5 Ablation Study
We evaluate the effectiveness of SPRLoss in Table.3. Compared with using the final answer correctness as process-level supervision for PRM training, our proposed self-supervised optimization method achieves larger performance gains on both 1.5B and 7B models. Furthermore, Fig.8 shows the prediction score gap between correct and incorrect solutions. Compared to the BCELoss, SPRLoss demonstrates stronger discriminative capability with a larger score gap. This indicates that treating final answer correctness as process-level labels introduces substantial label noise, which harms the optimization. In contrast, SPRLoss leverages self-supervised signals to reduce the impact of noisy supervision, leading to more stable and accurate training.
@@
5.6 Extend on MCTS
Since SPRM produces process scores for each step, our reflective generative models can be naturally used in step-level search-based TTS methods such as Monte Carlo Tree Search (MCTS). In our setup, instead of performing full simulations to the end of a reasoning trajectory, we directly use SPRM to estimate the value of each node during the search. This enables a more efficient evaluation at each expansion step. In the expanding stage, we expand 4 children for the selected node and generate 1024 tokens in each child node. To balance the computation cost, we set the maximum number of tokens in the MCTS process from 0 (without MCTS) to 160k for each question.
Table.4 presents the performance of MCTS on the AIME24 dataset. By increasing the maximum number of searching tokens, the performance improved from 39.3 to 52.8, demonstrating the effectiveness of SPRM in providing high-quality step-level guidance in the reasoning stage. However, the performance of MCTS is still lower than the results under the Best-of-N strategy reported in Table 1. This is primarily due to the computational overhead inherent in tree-based search methods, which leads to incomplete searching in our setting. Nonetheless, the observed gains over the baseline validate the potential of our reflective generative models for integration with more advanced search-based inference methods.
@@ 6 Conclusion
In this work, we propose a novel Reflective Generative Form, which enables a single LLM to both generate and select high-quality reasoning trajectories for Test-Time Scaling (TTS). Based on this form, we introduce our first reflective generative model, MetaStone-S1, which achieves comparable performance to the OpenAI o3-mini series. Specifically, we design a unified interface that integrates the policy model and process reward model (PRM) within a single network, resulting in low parameter overhead and efficient TTS inference. To further reduce the dependence on costly process-level annotations for PRM training, we present a self-supervised process reward model (SPRM) that learns process-level evaluation using only final answer annotations. MetaStone-S1, with 32B parameters, achieves strong performance across mathematics, coding, and Chinese reasoning benchmarks, outperforming a series of open-source and closed-source models. In addition, our experimental analyses of the aha moment and scaling law further demonstrate the effectiveness of our Reflective Generative Form. In future work, we plan to explore its capability for more efficient step-level search-based TTS for real-time reasoning enhancement.